OpenStudy (gabylovesyou):

6sqrt(v) - sqrt(4v^2) - sqrt(36v) + sqrt(v^2) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have to simplify each piece

jimthompson5910 (jim_thompson5910):

so for instance sqrt(4v^2) = sqrt((2v)^2) = 2v